# User Simulation for Evaluating Information Access Systems

### v1

@@ 4.2.2 Measures for Search Sessions
The measures discussed above are all restricted to evaluation of the results of a single query. However, retrieval tasks often require a user to reformulate queries, which can be due to an ineffective initial query or an exploratory information need. In such cases, it is more important to evaluate the overall performance of a whole search session. To this end, some measures defined on a single query have been generalized to measure the performance over a session that consists of multiple queries. For example, nDCG has been extended to session-based search, taking user effort implicitly into account, leading to the Session nDCG (sDCG) measure [Jarvelin:2008:ECIR]. The main idea is to discount the results returned from a later query in a session. Technically, the DCG discount vectors from all the queries can be concatenated by applying an additional discount coefficient based on the query position in the session, thus enabling the computation of nDCG for the whole session by simply pretending the concatenated results were a single ranked list of documents. From a user simulation perspective, the main difference from a single query-based measure is the generalization of the sequential browsing of a ranked list to sequential browsing of multiple ranked lists in a session. A more general formal framework for modeling the user browsing behaviour and computing an Expected Global Utility over a session has been proposed in [Yang:2009:ICTIR], which enables modeling the uncertainty of a user’s browsing behaviour and computes the expected utility w.r.t. the distribution of all possible user browsing behaviours. The application of the framework to derive a session-based nDCG measure is discussed in [Yang:2009:ICTIR]. Similar ideas have been explored in [Kanoulas:2011:SIGIR], where the idea of modeling a user’s browsing behaviour in a session using a “path” was introduced, which can be roughly interpreted as modeling the perceived ranking of all the documents a user has interacted with in a session as a single ranked list. Any measure can then be defined based on such a perceived ranked list for the whole session. Two strategies have been explored in depth for implementing the idea, leading to model-free session evaluation (no explicit modeling of user behaviour, using the best path among a set of possibilities) and model-based session evaluation (with assumptions of user’s behaviour, using the expectation according to the assumed user behaviour distribution).
@@
Compared with single query-based evaluation measures, those session-based evaluation measures can be regarded as evaluation using more realistic and more sophisticated user behaviour models (browsing through a whole session vs. through a single ranked list). However, the browsing behaviour has not been modeled in detail in such measures; for example, the clicking decision that drives the browsing behaviour has not been modeled in those measures, even though click models have been separately studied [Chuklin:2015:Book]. The Interactive IR Probability Ranking Principle (IIR-PRP) [Fuhr:2008:IRJ] theoretically integrates a user’s clicking decision with a measure of the overall utility of a ranked list with consideration of dependency of document utility in the list. While IIR-PRP was proposed for deriving an optimal ranking algorithm, it can also be regarded as defining an evaluation measure based on relevance gain and cost of examining documents, with explicit modeling of sequential browsing decisions. All these measures, however, are based on the assumption of sequential browsing, which is only applicable when the search results are displayed as a ranked list of results (or there is a natural way for a user to perform sequential interaction with the results). Modern search and recommender systems provide much more sophisticated user interfaces that can easily go beyond sequential browsing (e.g., results organized in a grid layout). In the next section, we will discuss a more general simulation-based evaluation framework for evaluating retrieval systems that support more sophisticated user interactions, including those going beyond sequential browsing.
@@ 4.3 A General Simulation-based Evaluation Framework
In [Zhang:2017:ICTIR], such a general evaluation framework is formally defined as follows. Let $S$ be a system, $U$ be a user, and $I$ be the whole process of the interaction of $U$ and $S$ to finish task $T$. We would be interested in measuring the system’s performance based on $I$. From a user’s perspective, we can measure the performance in two dimensions: (1) Interaction Reward, $R(I,T,U,S)$, which measures the total reward the user has received via the interaction, and (2) Interaction Cost, $C(I,T,U,S)$, which measures the total cost of the interaction. In general, the more interaction actions the user makes, the more reward the user can potentially receive and the more cost the user would have to bear (since the user needs to make more effort). If one single measure is needed, the reward and cost can be combined, which can be in many different forms. The need for modeling both reward and cost was already recognized in the Interactive IR Probability Ranking Principle [Fuhr:2008:IRJ], where the two are integrated in a decision model simulating how a user makes browsing decisions. In [Zhang:2017:ICTIR], the integration of the reward and cost is regarded as the foundation of an evaluation measure. Intuitively, this reflects the intuition that an ideal system should provide maximize reward to a user while minimizing the cost/effort required on the user.
@@
In [Zhang:2017:ICTIR], the general framework is further refined to assume that $I$ is a sequence of user actions taken in response to a sequence of Interface Card, generated by system $S$. Under such an assumption, the interaction sequence $I=((z_{1},a_{1},q_{1}),(z_{2},a_{2},q_{2}),...,(z_{n},a_{n},q_{n}))$, where $a_{i}$ is a user action, $q_{i}$ is an interface card (i.e., a dynamic user interface) generated by the system, and $z_{i}$ is a representation of the user’s state during the interaction. The meaning of $(z_{i},a_{i},q_{i})$ is that while in state $z_{i}$, the user has taken action $a_{i}$ (e.g., entering a query), which leads to the system generating an interface card $q_{i}$ (e.g., showing a list of search results). With this refinement, the reward and cost can be further decomposed into the action-level reward.
It is easy to see that such a general framework is applicable to evaluation of any interactive information access system, or even any assistive AI system. When the user actions are restricted to browsing of documents in a ranked list sequentially, the framework can naturally cover some specific frameworks such as C/W/L [Moffat:2017:TOIS], C/W/L/A [Moffat:2022:SIGIR] and the model-based measure framework [Carterette:2011:SIGIR] as special cases, thus also covering virtually all the existing specific IR measures.
Analysis of some existing measures in this general framework has revealed interesting interpretations of those measures, which may guide the choice of appropriate measures for a specific purpose of evaluation. For example, Average Precision can be interpreted as the expectation of a retrieval system’s utility for a set of users whose tasks have variable levels of recall [Zhang:2017:ICTIR].
It was shown in [Zhang:2017:ICTIR] that the general framework, with assistance from the Interface Card Model [Zhang:2015:SIGIR], can be used to evaluate an interactive IR system with a computationally generated dynamic browsing interface using user simulation. Specifically, the task was to browse into a target relevant document in a collection (known-item browsing), and the goal was to compare a system which can generate a browsing interface dynamically in adaptation to both the display size and the system’s confidence in the inference about the user’s interest [Zhang:2015:SIGIR] with a baseline system that has an adaptive interface to display size that can choose to use at any time. Simple user simulators were created and the system performance is measured based on the effort required in order to reach the target item (the number of clicks required). The results using user simulation were found to be consistent with those from using real user studies [Zhang:2017:ICTIR].
In general, with the simulation framework, we can potentially evaluate a search engine or a recommender system using any sophisticated user simulator. In the Chapters 6 and 7, we will discuss how such sophisticated user simulators can be constructed.
@@
Construction of user simulators requires specifying the functions of the system that a user interacts with as well as modeling how a user makes a decision between multiple actions that could be potentially taken by using each function offered by a system. As we intend to cover broadly multiple types of information access systems, including particularly search engines, recommender systems, and conversational assistants, we would need to address the question of what their functions are exactly. There are many possible answers to this question. In this book, we view search engines as tools that help people find information, recommender systems as tools that aid people in making better choices, and conversational assistants as tools that support users in completing their tasks via natural language interactions (cf. Section 1.1). In order to successfully simulate user behaviour when interacting with such systems, it is important to have a broad understanding of the cognitive processes involved, which play a crucial role in shaping users’ behaviour and decision-making. In this chapter, we first provide a high-level overview of research on how people seek information (Section 5.1) and discuss various conceptual models that can provide theoretical guidance for modeling processes and decisions from an individual’s perspective. We then discuss how users make choices in the context of recommender systems (Section 5.2). Finally, in Section 5.3, we discuss how to model decision-making processes mathematically using Markov decision processes (MDP). The MDP framework provides a general formal framework for constructing user simulators, which we will use to discuss specific user simulation techniques in the next two chapters, simulating interactions with search and recommender systems (Chapter 6) and interactions with conversational assistants (Chapter 7).
@@ 5.1 Conceptual Models of Information Seeking
The field of library and information science has had an interest in developing models of information behaviour based on a cognitive viewpoint since the late 1970s [deMey:1977:CC]. The central idea of the cognitive viewpoint is that information processing is human-oriented and thus draws attention to human perception, cognition, and structures of knowledge when interacting with information [Wilson:1984:SSIS]. This stood in stark contrast with the predominantly system-focused IR research of that time, centered around the representation and matching of queries and documents (cf. Fig. 5.1). Studies examining user’s search behaviour began in the 1980s, focusing on intermediary-assisted search taking place in libraries [Pejtersen:1980:TAIR, Ingwersen:1982:JD, Ellis:1989:JD]. The component of intermediary is later replaced by the construct of an interface/intermediary, allowing for interaction to occur between an user and an IR system [Ingwersen:1996:JD]. During early 1990s, a cognitive turn took place in information seeking and retrieval, from focusing on the technological aspects only to a more human-oriented approach to information interaction [Ingwersen:2005:book].
@@
Before we get into the specifics, a general note on the limitations and relevance of these models. As Wilson:1999:JD explains, “most models in the general field of information behaviour are […] statements, often in the form of diagrams, that attempt to describe an information-seeking activity, the causes and consequences of that activity, or the relationships among stages in information-seeking behaviour. […] The limitation of this kind of model, however, is that it does little more than provide a map of the area and draw attention to gaps in research.” Furthermore, the validation of information seeking models to large-scale behaviour data remains largely an open question [White:2016:book]. That said, these pioneering models are still relevant because they describe “relationships between information seekers, human intermediaries, information resources and information systems” [Savolainen:2018:JD]. Conceptual models can provide theoretical guidance for building user models and inform the high-level design of user simulators. These works also provide useful directions for future research to make simulators more realistic—for example, the representation of context and modeling of the user’s knowledge state have been largely unexplored to date (cf. Section 8.4.1).
@@ 5.1.1 Cognitive Models
Ingwersen:1996:JD specifies elements of a cognitive IR theory, combining both system-oriented and cognitive research. IR elements include the representation of information objects, queries, and retrieval techniques. Cognitive elements include the individual’s work task, cognitive state, problem space, and information behaviour, as well as their socio-organizational environment. The strength of the model lies in the detailed description of essential processes from both the user and system perspectives. On the other hand, it paints a fairly complicated picture of information interactions and cognitive structures, while the model’s applicability to system evaluation is limited [Savolainen:2018:JD]. Ingwersen:2005:book elaborate further on this model and present a generalized research framework for information seeking and retrieval (IS&R), illustrated in Fig. 5.2. It consists of five main components: information objects, information technology, interface, information seeker, and context. There are two possible types of relationships between the components: (1) cognitive transformation and influence over time, and (2) longitudinal interaction of cognitive structures. The IS&R model treats the information seeker as the central actor and emphasizes “the interaction between the information seeker(s) and the environment surrounding that individual, also over time” [Ingwersen:2005:book]. The model further emphasizes the mutual dependencies between the different cognitive components, including intra-component structures. While it represents the most systematic and holistic framework for information interaction to date, it remains at a very high level of conceptualization and does not describe the nature of interactions in greater detail [Savolainen:2018:JD].
@@ 5.1.3 Strategic Models
The berry-picking model, proposed by Bates:1989:OR, considers information seeking analogous to foragers looking for food. One central point of this model is that searchers’ needs are not satisfied by a single set of retrieved results (as assumed by the classic more of IR, cf. Fig. 5.1). Rather, the information they are looking for is scattered like berries on bushes, and needs to be collected bit by bit. Another main point of the model is that as searchers encounter new pieces of information and references along the way, those might give them new ideas and directions to follow. Thus, search is an evolving process and as it progresses, not just the queries but the underlying information needs are also continually shifting, in part or whole; see Fig. 5.4 for an illustration. The berry-picking model is supported by observational studies [ODay:1993:CHI, Borgman:1996:JASIST] and the assumption that information needs are satisfied as a result of multiple queries are also captured in many of the session- or task-based models in Web search [White:2010:SIGIR, Eickhoff:2014:WSDM].
Information foraging theory [Pirolli:1999:PR] applies ideas from optimal foraging theory to understand how people search for information and interact with potential sources of information. Optimal foraging theory explains how animals maximize their fitness while they search for food, i.e., gain the most energy for the lowest cost [Stephens:1986:book]. Information foraging theory thus assumes that humans use the same innate foraging mechanisms and behave in a way to maximize their rate of gaining valuable information over time. Central to the theory is the notion of a patch—in the wild, this is an area where food can be acquired. Foragers need to decide how long they want to stay in a patch before moving to the next patch. Applied to the task of information seeking, staying on a patch corresponds to examining the results presented on a SERP, while moving between patches is akin to issuing a new query [Azzopardi:2015:ICTIR]. Another central concept of information foraging theory is information scent—similar to how animals can rely on scents to indicate their chances of finding prey, searchers can examine cues presented to them on web pages or SERPs to help locate information [Woodruff:2001:CHI, Olston:2003:TCHI, Pirolli:2007:book]. There is empirical evidence showing that when information scent starts to decrease, searchers transition to other information sources [Card:2001:CHI] and that users examine the SERP to greater depths if it appears to contain many relevant results [Card:2001:CHI, Wu:2014:SIGIR].

### v2

@@ 4.2.2 Measures for Search Sessions
The measures discussed above are all restricted to evaluation of the results of a single query. However, retrieval tasks often require a user to reformulate queries, which can be due to an ineffective initial query or an exploratory information need. In such cases, it is more important to evaluate the overall performance of a whole search session. To this end, some measures defined on a single query have been generalized to measure the performance over a session that consists of multiple queries. For example, NDCG has been extended to session-based search, taking user effort implicitly into account, leading to the Session NDCG (SDCG) measure [Jarvelin:2008:ECIR]. The main idea is to discount the results returned from a later query in a session. Technically, the DCG discount vectors from all the queries can be concatenated by applying an additional discount coefficient based on the query position in the session, thus enabling the computation of NDCG for the whole session by simply pretending the concatenated results were a single ranked list of documents. From a user simulation perspective, the main difference from a single query-based measure is the generalization of the sequential browsing of a ranked list to sequential browsing of multiple ranked lists in a session. A more general formal framework for modeling the user browsing behaviour and computing an Expected Global Utility over a session has been proposed in [Yang:2009:ICTIR], which enables modeling the uncertainty of a user’s browsing behaviour and computes the expected utility w.r.t. the distribution of all possible user browsing behaviours. The application of the framework to derive a session-based NDCG measure is discussed in [Yang:2009:ICTIR]. Similar ideas have been explored in [Kanoulas:2011:SIGIR], where the idea of modeling a user’s browsing behaviour in a session using a “path” was introduced, which can be roughly interpreted as modeling the perceived ranking of all the documents a user has interacted with in a session as a single ranked list. Any measure can then be defined based on such a perceived ranked list for the whole session. Two strategies have been explored in depth for implementing the idea, leading to model-free session evaluation (no explicit modeling of user behaviour, using the best path among a set of possibilities) and model-based session evaluation (with assumptions of user behaviour, using the expectation according to the assumed user behaviour distribution).
@@
Compared with single query-based evaluation measures, those session-based evaluation measures can be regarded as evaluation using more realistic and more sophisticated user behaviour models (browsing through a whole session vs. through a single ranked list). However, the browsing behaviour has not been modeled in detail in such measures; for example, the clicking decision that drives the browsing behaviour has not been considered, even though click models have been separately studied [Chuklin:2015:Book]. The Interactive IR Probability Ranking Principle (IIR-PRP) [Fuhr:2008:IRJ] theoretically integrates a user’s clicking decision with a measure of the overall utility of a ranked list with consideration of dependency of document utility in the list. While IIR-PRP was proposed for deriving an optimal ranking algorithm, it can also be regarded as defining an evaluation measure based on relevance gain and cost of examining documents, with explicit modeling of sequential browsing decisions. All these measures, however, are based on the assumption of sequential browsing, which is only applicable when the search results are displayed as a ranked list (or there is a natural way for a user to perform sequential interaction with the results). Modern search and recommender systems provide much more sophisticated user interfaces that can easily go beyond sequential browsing (e.g., results organized in slates). In the next section, we will discuss a more general simulation-based evaluation framework for evaluating retrieval systems that support more sophisticated user interactions, including those going beyond sequential browsing.
@@ 4.3 A General Simulation-Based Evaluation Framework
In [Zhang:2017:ICTIR], such a general evaluation framework is formally defined as follows. Let $S$ be a system, $U$ be a user, and $I$ be the whole process of the interaction of $U$ and $S$ to finish task $T$. We would be interested in measuring the system’s performance based on $I$. From a user’s perspective, we can measure the performance in two dimensions: (1) Interaction Reward, $R(I,T,U,S)$, which measures the total reward the user has received via the interaction, and (2) Interaction Cost, $C(I,T,U,S)$, which measures the total cost of the interaction. In general, the more interaction actions the user makes, the more reward the user can potentially receive and the more cost the user would have to bear (since the user needs to make more effort). If one single measure is needed, the reward and cost can be combined, which can be in many different forms. The need for modeling both reward and cost was already recognized in the Interactive IR Probability Ranking Principle [Fuhr:2008:IRJ], where the two are integrated in a decision model simulating how a user makes browsing decisions. In [Zhang:2017:ICTIR], the integration of the reward and cost is regarded as the foundation of an evaluation measure. This reflects the intuition that an ideal system should maximize reward to a user while minimizing the cost or effort required of the user. Besides these general frameworks, there are many specific models, notably economic IR models, where cost and gain are explicitly modeled (see, e.g.,  [Azzopardi:2011:SIGIR] and many follow-up studies).
@@
In [Zhang:2017:ICTIR], the general framework is further refined to assume that $I$ is a sequence of user actions taken in response to a sequence of interface cards, generated by system $S$. Under such an assumption, the interaction sequence is $I=((z_{1},a_{1},q_{1}),(z_{2},a_{2},q_{2}),...,(z_{n},a_{n},q_{n}))$, where $a_{i}$ is a user action, $q_{i}$ is an interface card (i.e., a dynamic user interface) generated by the system, and $z_{i}$ is a representation of the user’s state during the interaction. The meaning of $(z_{i},a_{i},q_{i})$ is that while in state $z_{i}$, the user has taken action $a_{i}$ (e.g., entering a query), which leads to the system generating an interface card $q_{i}$ (e.g., showing a list of search results). With this refinement, reward and cost can be further decomposed into action-level reward.
It is easy to see that this general framework is applicable to the evaluation of any interactive information access system, or even any assistive AI system. When the user actions are restricted to browsing of documents in a ranked list sequentially, the framework can naturally cover some specific frameworks such as C/W/L [Moffat:2017:TOIS], C/W/L/A [Moffat:2022:SIGIR] and the model-based measure framework [Carterette:2011:SIGIR] as special cases, thus also covering virtually all the existing specific IR measures.
The analysis of existing measures within this general framework has revealed interesting interpretations, potentially guiding the choice of appropriate measures for specific evaluation purposes. For example, Average Precision can be interpreted as the expectation of a retrieval system’s utility for a set of users whose tasks have variable levels of recall [Zhang:2017:ICTIR].
It was shown in [Zhang:2017:ICTIR] that the general framework, with assistance from the Interface Card Model [Zhang:2015:SIGIR], can be used to evaluate an interactive IR system with a computationally generated dynamic browsing interface using user simulation. Specifically, the task was to browse to a target relevant document in a collection (known-item browsing), and the goal was to compare a system which can generate a browsing interface dynamically in adaptation to both the display size and the system’s confidence in the inference about the user’s interest [Zhang:2015:SIGIR] with a baseline system that has an adaptive interface to display size that can choose to use at any time. Simple user simulators were created and the system performance is measured based on the effort required in order to reach the target item (the number of clicks required). The results using user simulation were found to be consistent with those from using real user studies [Zhang:2017:ICTIR].
In general, with the simulation framework, we can potentially evaluate a search engine or a recommender system using any sophisticated user simulator. In Chapters 6 and 7, we will discuss how such sophisticated user simulators can be constructed.
@@
Construction of user simulators requires specifying the functions of the system that a user interacts with as well as modeling how a user makes a decision between multiple actions that could be potentially taken based on the functionality of the system. As we intend to cover broadly multiple types of information access systems, including particularly search engines, recommender systems, and conversational assistants, we would need to address the question of what exactly their functions are. There are many possible answers to this question. In this book, we view search engines as tools that help people find information, recommender systems as tools that aid people in making better choices, and conversational assistants as tools that support users in completing their tasks via natural language interactions (cf. Section 1.1). In order to successfully simulate user behaviour when interacting with such systems, it is important to have a broad understanding of the cognitive processes involved, which play a crucial role in shaping users’ behaviour and decision-making. In this chapter, we first provide a high-level overview of research on how people seek information (Section 5.1) and discuss various conceptual models that can provide theoretical guidance for modeling processes and decisions from an individual’s perspective. We then discuss how users make choices in the context of recommender systems (Section 5.2). Finally, in Section 5.3, we discuss how to model decision-making processes mathematically using Markov decision processes (MDP). The MDP framework provides a general formal framework for constructing user simulators, which we will use to discuss specific user simulation techniques in the next two chapters, simulating interactions with search and recommender systems (Chapter 6) and interactions with conversational assistants (Chapter 7), respectively.
@@ 5.1 Conceptual Models of Information Seeking
The field of library and information science has had an interest in developing models of information behaviour based on a cognitive viewpoint since the late 1970s [deMey:1977:CC]. The central idea of the cognitive viewpoint is that information processing is human-oriented and thus draws attention to human perception, cognition, and structures of knowledge when interacting with information [Wilson:1984:SSIS]. This stood in stark contrast with the predominantly system-focused IR research of that time, centered around the representation and matching of queries and documents (cf. Fig. 5.1). Studies examining user’s search behaviour began in the 1980s, focusing on intermediary-assisted search taking place in libraries [Pejtersen:1980:TAIR, Belkin:1982:JD, Ingwersen:1982:JD, Ellis:1989:JD]. The component of intermediary is later replaced by the construct of an interface/intermediary, allowing for interaction to occur between a user and an IR system [Ingwersen:1996:JD]. During early 1990s, a cognitive turn took place in information seeking and retrieval, from focusing on the technological aspects only to a more human-oriented approach to information interaction [Ingwersen:2005:book].
@@
Before we get into the specifics, a general note on the limitations and relevance of these models. As Wilson:1999:JD explains, “most models in the general field of information behaviour are […] statements, often in the form of diagrams, that attempt to describe an information-seeking activity, the causes and consequences of that activity, or the relationships among stages in information-seeking behaviour. […] The limitation of this kind of model, however, is that it does little more than provide a map of the area and draw attention to gaps in research.” Furthermore, the validation of information seeking models based on large-scale behaviour data remains largely an open question [White:2016:book]. That said, these pioneering models are still relevant because they describe “relationships between information seekers, human intermediaries, information resources and information systems” [Savolainen:2018:JD]. Conceptual models can provide theoretical guidance for building user models and inform the high-level design of user simulators. These works also provide useful directions for future research to make simulators more realistic (cf. Section 10.4.1). For example, the representation of context and modeling of the user’s knowledge state have been largely unexplored to date—this gap presents opportunities to link to research on “searching as learning,” where understanding how users acquire knowledge during search sessions is crucial [Vakkari:2016:JIS, Ghosh:2018:CHIIR].
@@ 5.1.1 Cognitive Models
Ingwersen:1996:JD specifies elements of a cognitive IR theory, combining both system-oriented and cognitive research. IR elements include the representation of information objects, queries, and retrieval techniques. Cognitive elements include the individual’s work task, cognitive state, problem space, and information behaviour, as well as their socio-organizational environment. The strength of the model lies in the detailed description of essential processes from both the user and system perspectives. On the other hand, it paints a fairly complicated picture of information interactions and cognitive structures, while the model’s applicability to system evaluation is limited [Savolainen:2018:JD]. Ingwersen:2005:book elaborate further on this model and present a generalized research framework for information seeking and retrieval (IS&R), illustrated in Fig. 5.2. It consists of five main components: information objects, information technology, interface, information seeker, and context. There are two possible types of relationships between the components: (1) cognitive transformation and influence over time, and (2) longitudinal interaction of cognitive structures. The IS&R model treats the information seeker as the central actor and emphasizes “the interaction between the information seeker(s) and the environment surrounding that individual, also over time” [Ingwersen:2005:book]. The model further emphasizes the mutual dependencies between the different cognitive components, including intra-component structures. While it represents the most systematic and holistic framework for information interaction to date, it remains at a very high level of conceptualization and does not describe the nature of interactions in greater detail [Savolainen:2018:JD]. Such a framework, however, provides useful guidance for the design of realistic user simulators.
@@ 5.1.3 Strategic Models
The berry-picking model, proposed by Bates:1989:OR, considers information seeking analogous to foragers looking for food. One central point of the model is that searchers’ needs are not satisfied by a single set of retrieved results (as assumed by the classic more of IR, cf. Fig. 5.1). Rather, the information they seek is scattered like berries on bushes, and needs to be collected bit by bit. Another main point of the model is that as searchers encounter new pieces of information and references along the way, those might give them new ideas and directions to follow. Thus, search is an evolving process and as it progresses, not just the queries but the underlying information needs are also continually shifting, in part or whole; see Fig. 5.4 for an illustration. The berry-picking model is supported by observational studies [ODay:1993:CHI, Borgman:1996:JASIST] and the assumption that information needs are satisfied as a result of multiple queries are also captured in many of the session- or task-based models in Web search [White:2010:SIGIR, Eickhoff:2014:WSDM].
Information foraging theory [Pirolli:1999:PR] applies ideas from optimal foraging theory to understand how people search for information and interact with potential sources of information. Optimal foraging theory explains how animals maximize their fitness while they search for food, i.e., gain the most energy for the lowest cost [Stephens:1986:book]. Information foraging theory thus assumes that humans use the same innate foraging mechanisms and behave in a way to maximize their rate of gaining valuable information over time. Central to the theory is the notion of a patch—in the wild, this is an area where food can be acquired. Foragers need to decide how long they want to stay in a patch before moving to the next patch. Applied to the task of information seeking, staying on a patch corresponds to examining the results presented on a SERP, while moving between patches is akin to issuing a new query [Azzopardi:2015:ICTIR]. Another central concept of information foraging theory is information scent—similar to how animals can rely on scents to indicate their chances of finding prey, searchers can examine cues presented to them on web pages or SERPs to help locate information [Woodruff:2001:CHI, Olston:2003:TCHI, Pirolli:2007:book]. There is empirical evidence showing that when information scent starts to decrease, searchers transition to other information sources [Card:2001:CHI] and that users examine the SERP to greater depths if it appears to contain many relevant results [Card:2001:CHI, Wu:2014:SIGIR]. Azzopardi:2018:SIGIR’s information foraging metric [Azzopardi:2018:SIGIR] is a useful example of Pirolli:1999:PR’s work translated directly into a user model and a metric.
@@
It should also be noted that all three kinds of models discussed in this section are interconnected, each offering a unique perspective.